1. Suppose the price of CANNIBAL WEED (CW) Inc. closes at $69 on October 16, 2018 after the announcement that an extraordinary dividend of $3 per share will be distributed on October 17, 2018 before trading starts.
(a) Based on this information, what price would you expect the stock to trade at on October 17 after the dividend has been distributed?
(b) Based on your answer to part (a), what is the return for an investor that buys the stock just before the closing on October 16 and sells it again immediately after trading starts on October 17? Explain your answer.
(c) Suppose the opening price for the stock on October 17, the day cannabis has been legalized, is $64. What does this imply for the ecient market hypothesis?
Answer #1

(a) Expected price = Ex dividend price = Pre dividend price - dividend = 69 - 3 = $ 66

(b) P0 = Buy price = 69; Dividend, D = 3; Sale Price, P1 = 66; Hence the return = (P1 + D - P0) / P0 = (66 + 3 - 69) / 69 = 0%

Thus in the ideal scenario, dividend doesn't bring any additional value for the shareholders. The price of the stock falls to the extent of dividend distributed to remove any abnormality or arbitrage in the market.

(c) Price should have been $ 66 on account of dividend distribution as calculated in part (a). Since legalization of cannabis is expected to work in favor of CW, the stock price should have increased. Hence, the price should have been definitely $ 66 or more. However it opened at a lower price. This simply implies the lack of strong or even semi strong form of efficiency. At the best we can say that market has a weak for of efficiency.